consider a home mortgage of 225000 at a fixed apr 4.5% for 25 years... I need all

ai2 Me·4.D-S Calculate loan payments, tota This Question: 1 pt 1 of Consider a home mortgage of $225,000 at a fixed APR of 4 5% for 25 years. a. Calculate the monthly payment b. Determine the total amount paid over the term of the loan. c. Of the total amount paid, what percentage is paid toward the principal and what percentage is paid for interest a. The monthly payment is $ 1,218.91 (Do not round until the final answer. Then round to the nearest cent as needed.) b. The total amount paid over the term of the loan is $ 365,673 22 (Round to the nearest cent as needed.) c of the total amount paid □96 paid toward the principal, and % is paid for interest (Round to one decimal place as needed.)
Answer #1

1.
PV=225000
N=25*12=300
I/Y=4.25%/12
FV=0
CPT PMT=1218.91

2.
Total amount paid over the term=1218.91*12*25=365673

3.
Principal Paid=225000
Interest paid=365673-225000=140673
% paid toward the principal=225000/365673=61.53038% and 100-61.53038=38.46962% is paid toward interest